OpenStudy (chacha1234):

Which system of equations below has no solution? A. y = 4x + 5 and y = 4x – 5 B. y = 4x + 5 and 2y = 8x + 10 C. y = 4x + 5 and y = 1/3 + 5 D. y = 4x + 5 and y = 8x + 10

OpenStudy (photon336):

so what you could do is set each pair of equations equal to each-other and if there is no solution you'll get something like 2 =3 when you plug one value into the other equation. or something like that. try solving the first one by setting both equations equal to each-other @ChaCha1234 start with the first one

OpenStudy (chacha1234):

I think it's the first one?

OpenStudy (chacha1234):

Ya it's the first one!

OpenStudy (chacha1234):

Am I right?

OpenStudy (photon336):

show me why it is the first one

OpenStudy (photon336):

exactly no solution.

OpenStudy (chacha1234):

Can you help with one more?

OpenStudy (photon336):

5 = -5

OpenStudy (photon336):

yeah tag me.

OpenStudy (chacha1234):

@Photon336 A system of equations is given below. y=1/2x-3 and y=-1/2x-3 Which of the following statements best describes the two lines? They have the same slope but different y-intercepts, so they have no solution. They have the same slope but different y-intercepts, so they have one solution. They have different slopes but the same y-intercept, so they have no solution. They have different slopes but the same y-intercept, so they have one solution.

OpenStudy (photon336):

y = mx+b where b = the y intercept and m= the slope

OpenStudy (photon336):

you could try setting the equal to each-other and trying solve them. 0.5x-3 = -0.5x-3 x = -3+3 x = 0 0, -3 they have one solution,

OpenStudy (chacha1234):

So.. B @Photon336 ?

OpenStudy (photon336):

they have a different slope, same y intercept, but one solution.